OpenStudy (anonymous):

solve for x x-3/4 + x+3/7 =3

OpenStudy (anonymous):

well what r the directions

OpenStudy (anonymous):

o sollve 4 x haha

OpenStudy (anonymous):

lol!!! duh!!!!

OpenStudy (anonymous):

well 1st u have 2 get the variables alone

OpenStudy (anonymous):

summer finks....

OpenStudy (anonymous):

x-3/4 + x+3/7 =3 |Original question x+x-3/4+3/7=3 |Move x next to each other for organization 2x-3/4+3/7=3 |Add the x together 2x-21/28+12/28=84/28 |convert 3/4 and 3/7 to same denominator 2x-9/28=84/28 |Add two fractions 2x=93/28 | Add -9/28 to both sides x=93/56 | Divide by two to get this answer :D

OpenStudy (anonymous):

summer......wats the answer???

OpenStudy (anonymous):

wow confusing there topsecret

OpenStudy (anonymous):

well 1st get the variables alone

OpenStudy (anonymous):

summer, go 2 my new question!!!!!!

OpenStudy (anonymous):

ok